# FLoRA: An Advanced AI-Powered Engine to Facilitate Hybrid Human-AI Regulated Learning

### v2

@@ 5.3 Implications for Future Design and Research
Second, providing learners with direct access to visual analytics presents a promising avenue for future development. Learner-facing dashboards that depict real-time SRL activities can enhance metacognitive monitoring, prompting students to reflect and make targeted adjustments to their learning strategies (Authors, 2023). Such tools can help cultivate meta-level awareness and encourage the gradual shift from external scaffold reliance to self-directed regulation.
Third, integrating Bloom’s taxonomy into FLoRA ’s writing analytics demonstrated its capacity to improve the accuracy of students’ self-assessment and metacognitive monitoring (Tang et al., 2024). However, effective use of formative feedback based on cognitive domains requires preliminary learner training (Forehand, 2010). Expanding FLoRA to include educator training modules, which complement learner-focused SRL interventions with teacher-led scaffolding, could further enhance the platform’s ability to support a diverse user base.
Finally, as highlighted in Case 3, adaptive scaffolding should consider both learners’ SRL processes and their individual characteristics, including domain knowledge and feedback literacy. Future iterations of FLoRA should leverage GenAI’s capacity to dynamically integrate real-time learner conditions, potentially using multimodal data such as eye-tracking for attention and facial expressions for confusion. This level of personalization surpasses traditional rule-based systems and supports recent theoretical calls for holistic, learner-centered analytics frameworks (Cukurova, 2024; Molenaar, 2022b), ultimately advancing hybrid human–AI regulated learning environments.
@@ 5.4 Limitations of the Study and Future Directions
While the results demonstrate FLoRA ’s capacity to support hybrid Human-AI regulation and advance SRL-focused educational research, this study has certain limitations. First, the reliance on digital trace data restricts visibility into higher-order cognitive and emotional SRL processes such as self-motivation, frustration management, or contextual sense-making (Saint et al., 2020a). Incorporating multi-modal data sources, such as physiological signals or speech analysis, may enhance the comprehensiveness of SRL detection, especially motivational and affective aspects (Raković et al., 2024), in future development and research of FLoRA . Second, the experimental tasks varied significantly in domain and structure (e.g., reflective writing versus clinical reasoning), which may partially confound comparisons of scaffolding efficacy across contexts. Future studies could standardise task designs to isolate the effect of scaffolding interventions under more controlled conditions. Furthermore, longitudinal studies are needed to examine the lasting impact of hybrid Human-AI regulation on learners’ metacognitive engagement and SRL skill development over time. Finally, the ethical implications of embedding GenAI systems into personalised learning contexts remain under-explored. For example, anonymity, privacy concerns, and data use transparency should be more systematically addressed in both the research design and platform development stages to ensure compliance with ethical standards in diverse regions and populations.
@@ 6 Conclusion
The enhanced FLoRA Engine represents a significant advancement in supporting Hybrid Human-AI Regulated Learning (HHAIRL) by providing both powerful research tools and a flexible infrastructure for further development. By integrating GenAI-driven instrumentation tools, a real-time trace parser, and adaptive scaffolding, FLoRA enables a comprehensive exploration of SRL processes and the nuanced interactions between learners and AI. The system’s ability to capture, analyse, and visualize real-time data facilitates comparisons of learning patterns with and without AI involvement, offering deeper insights into the role of AI in SRL.
Furthermore, to foster collaboration and open research, FLoRA is made available under an open-source license, allowing educators and researchers to customise, extend, and adapt its functionalities to diverse educational contexts. This flexibility ensures that FLoRA not only enhances AI-supported SRL interventions but also contributes to the ongoing advancement of HHAIRL research. By bridging the gap between automated AI-driven support and human agency, FLoRA empowers learners to actively regulate their learning while enabling researchers to refine and expand our understanding of effective human-AI collaboration in education.
Appendix A Other Features
@@ A.1 Other Tools
Figure 15: Math tool
A.2 Admin Portal
The enhanced FLoRA Engine now features an upgraded Admin portal that encompasses comprehensive data management, data visualization, and instrumentation tools configuration functionalities. In the previous version, the log portal was limited to basic data access capabilities, which proved insufficient given the advancements in our instrumentation tools. With the current version capturing richer and more detailed data, solely supporting data downloads no longer meets the evolving needs of researchers.
Figure 16: Admin portal - data management
To address this limitation,the data management portal’s functionalities (displayed in Figure 16) have been significantly enhanced. Researchers can now not only access and download data but also view essential metadata such as the total number of participants in a given experiment, the average number of trace data records per participant, and the number of data entries generated by different tools. This additional information provides researchers with a clearer overview of their datasets, facilitating more informed analyses. The data search process has also been improved to support fuzzy matching using account names or experiment names. This feature streamlines the data retrieval process, making it more efficient for researchers to locate specific datasets relevant to their studies.
Figure 17: Admin portal in FLoRA – visual data analytics for researchers
In terms of visual learning analytics, FLoRA has various graphical representations to depict data from multiple perspectives. For instance, Figure 17 illustrates the proportion of time different students spend on various learning actions within a single task. Such visualizations enable researchers to gain deeper insights into students’ learning behaviours and patterns, thereby supporting more nuanced interpretations of the data.
Figure 18: Admin portal in in FLoRA – configuration of tools for different study settings
Finally, to enhance the usability of the FLoRA Engine for researchers, FLoRA now has instrumentation tools configuration utility (as shown in Figure 18). This tool assists researchers in selecting and setting up the specific tools they wish to employ in their studies. By simplifying the customization process, we aim to empower researchers to tailor the platform to their unique research needs effectively.
Appendix B Architecture
The enhanced architecture of the FLoRA Engine is illustrated in Figure 19. In this upgraded version, the FLoRA Engine adopts a microservice architecture, wherein each service is responsible for specific tasks, promoting modularity and scalability. As shown in Figure 19, the entire system is segmented into four primary components: the Instrumentation Tools, the Control Service, the OmniAI Service, and the Admin Portal.
Figure 19: FLoRA Architecture
The Instrumentation Tools, developed using JavaScript, serve as the frontend interface and are designed to integrate seamlessly into learning environments such as Moodle. These tools capture user interactions and learning behaviours, collecting a different of data streams including mouse movements, clicks, and interactions with various areas or tools. This data reflects learners’ engagement and the specific learning actions they undertake during the learning process.
Once collected, the data is transmitted to the Control Service via middleware technologies including Kafka, Redis, ElasticSearch, which facilitate efficient data streaming and handling. The Control Service persists the raw data into databases and performs automatic analyses to assess learners’ progress. For instance, it can identify learners’ engagement in SRL processes. Based on the analysis results, the system provides personalised support, such as scaffolding, to enhance the learning experience and address individual learner needs.
The Control Service is built upon the robust Java frameworks Spring Boot and Spring Cloud. These frameworks are widely recognised for facilitating the development of scalable, secure, and efficient microservices. Spring Boot streamlines the creation of production-ready applications by minimizing boilerplate code and configuration efforts, allowing for rapid development and deployment. Spring Cloud extends these capabilities by providing tools for distributed systems, such as service discovery, configuration management, and circuit breakers. The combined use of Spring Boot and Spring Cloud enhances the Control Service’s ability to handle large-scale user access and high concurrency, ensuring reliability and performance. Notably, industry-leading companies like Netflix and Alibaba utilise these frameworks to support their microservices architectures, handling vast amounts of data and client requests efficiently.
During students’ learning activities, the OmniAI Service manages all machine learning-related processing. This includes sending requests to various large language model (LLM) APIs, such as OpenAI’s ChatGPT and Google’s language models, as well as conducting text analyses using machine learning models like Sentence Transformers. The OmniAI Service performs these advanced computational tasks to provide intelligent responses, feedback, and analyses that enhance the learning experience. Developed using the Python Flask framework, the OmniAI Service benefits from Flask’s lightweight yet powerful features. Flask is well-suited for building microservices and handling machine learning tasks due to its simplicity and flexibility. Its minimalistic design allows developers to implement necessary functionalities without unnecessary overhead, while its extensive ecosystem and strong community support facilitate the incorporation of various extensions and tools. By utilizing Flask, the OmniAI Service efficiently manages the computational demands of machine learning operations while maintaining scalability and performance.
The Admin Portal is responsible for overseeing data management, data visualization, and the configuration of instrumentation tools, with backend processing managed by the Control Service. In previous versions, the log portal’s functionality was limited to basic data access, primarily allowing users to download data. With significant enhancements in the current version’s instrumentation tools, the data captured has become richer and more complex. As a result, merely supporting data downloads is no longer sufficient to meet researchers’ needs. Therefore, we have significantly enhanced the functionality of the data management portal (displayed in Figure 16).
The upgraded data management portal enables researchers not only to access and download data but also to view essential metadata and statistics. Researchers can now examine basic information such as the total number of participants in a study, the average number of trace data entries per participant, and the number of data entries generated by different tools. This additional information provides valuable insights into the data’s scope and composition, facilitating more informed analyses. To streamline the data retrieval process, the portal supports fuzzy matching in data searches using account names or experiment titles, enhancing usability and saving time.
For data visualization, the portal offers various graphical representations to visualize data from multiple perspectives. For instance, Figure 17 illustrates the proportion of time different students spend on various learning actions within a task. Such visualizations enable researchers to gain deeper insights into students’ learning behaviours and patterns, supporting more nuanced interpretations of the data. This enhanced visualization capability is crucial for identifying trends and informing instructional strategies.
To facilitate easier use of FLoRA by researchers, FLoRA has an instrumentation tools configuration utility (as shown in Figure 18). This utility assists researchers in selecting and configuring the specific tools they wish to employ in their studies, simplifying the customization process. By making the configuration of instrumentation tools more accessible, we empower researchers to tailor the platform to their unique research needs effectively.
The Admin Portal is built using React, a popular JavaScript library for building user interfaces. React’s component-based architecture allows for the creation of reusable UI components, leading to efficient development and maintenance. Its virtual DOM implementation enhances performance by minimizing direct manipulation of the actual DOM, resulting in faster rendering and a smoother user experience. Moreover, React facilitates the adoption of a front-end and back-end separation architecture (also known as a decoupled or headless architecture). This approach separates the user interface (front-end) from the server-side logic and data management (back-end), offering several benefits.
The separation of front-end and back-end concerns improves scalability, as each layer can be scaled independently to accommodate increased user loads or data volumes. It enhances flexibility and maintainability, allowing developers to update the front-end without affecting back-end services and vice versa. For example, if the number of users increases significantly, additional server instances can be allocated to the Control Service to handle increased data processing, while the front-end can be optimised separately to manage higher user traffic. This modularity is essential for maintaining performance and reliability in a system designed to support large-scale educational environments.
All middleware used in the FLoRA Engine, including Redis, ElasticSearch, and Kafka, are capable of cluster deployment to handle high-throughput data processing. Redis provides fast in-memory data storage, improving the efficiency of data retrieval operations. ElasticSearch offers powerful search and analytics capabilities, enabling quick querying of large datasets. Kafka facilitates real-time data streaming and messaging, ensuring efficient communication between services. By leveraging these robust middleware solutions, the FLoRA Engine effectively manages the substantial amounts of data generated by users, maintaining performance and reliability even under heavy loads.
@@ Appendix B Architecture
Appendix C GPT-based Scaffolding Triggering Logic
@@
Table 2: Learners’ Learning Conditions (Authors, 2025b)
Declaration of Competing Interest

### v3

@@ 5.3 Implications for Future Design and Research
Firstly, providing learners with direct access to visual analytics presents a promising avenue for future development. Learner-facing dashboards that depict real-time SRL activities can enhance metacognitive monitoring, prompting students to reflect and make targeted adjustments to their learning strategies (Li et al., 2023). Such tools can help cultivate meta-level awareness and encourage the gradual shift from external scaffold reliance to self-directed regulation.
Secondly, integrating Bloom’s taxonomy into FLoRA ’s writing analytics demonstrated its capacity to improve the accuracy of students’ self-assessment and metacognitive monitoring (Tang et al., 2024). However, effective use of formative feedback based on cognitive domains requires preliminary learner training (Forehand, 2010). Expanding FLoRA to include educator training modules, which complement learner-focused SRL interventions with teacher-led scaffolding, could further enhance the platform’s ability to support a diverse user base.
Finally, as highlighted in Case 3, adaptive scaffolding should consider both learners’ SRL processes and their individual characteristics, including domain knowledge and feedback literacy. Future iterations of FLoRA should leverage GenAI’s capacity to dynamically integrate real-time learner conditions, potentially using multimodal data such as eye-tracking for attention and facial expressions for confusion. This level of personalization surpasses traditional rule-based systems and supports recent theoretical calls for holistic, learner-centred analytics frameworks (Cukurova, 2024; Molenaar, 2022b), ultimately advancing hybrid human–AI regulated learning environments.
@@ 5.4 Limitations of the Study and Future Directions
While the results demonstrate FLoRA ’s capacity to support hybrid Human-AI regulation and advance SRL-focused educational research, this study has certain limitations. First, the reliance on digital trace data restricts visibility into higher-order cognitive and emotional SRL processes such as self-motivation, frustration management, or contextual sense-making (Saint et al., 2020a). Incorporating multi-modal data sources, such as physiological signals or speech analysis, may enhance the comprehensiveness of SRL detection, especially motivational and affective aspects (Raković et al., 2024), in future development and research of FLoRA . Second, the experimental tasks varied significantly in domain and structure (e.g., reflective writing versus clinical reasoning), which may partially confound comparisons of scaffolding efficacy across contexts. Future studies could standardise task designs to isolate the effect of scaffolding interventions under more controlled conditions. Furthermore, longitudinal studies are needed to examine the lasting impact of hybrid Human-AI regulation on learners’ metacognitive engagement and SRL skill development over time.
The ethical implications of embedding GenAI systems into personalised learning contexts remain under-explored. For example, anonymity, privacy concerns, and data use transparency should be more systematically addressed in both the research design and platform development stages to ensure compliance with ethical standards in diverse regions and populations. In addition, safeguards such as privacy-preserving data processing and informed consent mechanisms should be embedded into the platform design to ensure transparency and learner trust. Beyond ethics, the risk of learner over-reliance on AI support warrants attention: novice learners may risk becoming dependent on external scaffolds, while more advanced learners may benefit from complementary support (Koedinger and Aleven, 2007; Belland et al., 2017). Future research should therefore investigate how to calibrate scaffold intensity and adapt interventions to learner needs, balancing guidance with opportunities for autonomy and fostering long-term SRL skill development (Molenaar, 2022a).
@@ 6 Conclusion
The enhanced FLoRA Engine represents a significant step forward in supporting Hybrid Human-AI Regulated Learning by providing robust research tools and a flexible, open-source infrastructure for further development. Through the integration of GenAI-driven instrumentation tools, a real-time trace parser, and adaptive scaffolding, FLoRA enables detailed exploration of SRL processes and the complex dynamics between learners and AI. Its ability to capture and visualize real-time data facilitates the comparison of learning patterns with and without AI involvement, yielding valuable insights into the potential and limitations of AI in SRL.
However, it is important to acknowledge that the implementation of GenAI tools, while promising for enhancing immediate learner performance, is not without risks. If employed without careful pedagogical guidance, such interventions may inadvertently lessen learners’ metacognitive engagement and autonomy (Darvishi et al., 2024). Therefore, the use of FLoRA must be informed by explicit instructional strategies that promote reflective and critical engagement with AI, rather than reliance on automated support alone.
From a practical standpoint, FLoRA offers significant implications for various stakeholders. Educators can leverage FLoRA to design targeted SRL interventions and to monitor learner progress with greater granularity. Educational technology developers may benefit from its open-source framework, which allows for customization and integration with diverse learning environments. Institutional leaders, meanwhile, can utilize the system’s analytics to inform policy decisions and resource allocation in support of evidence-based, AI-enhanced education.
Looking ahead, several avenues for future research and development merit consideration. Expanding FLoRA ’s capabilities to incorporate multimodal data—such as voice, gesture, or physiological signals—could yield a more comprehensive understanding of SRL processes and learner-AI interactions. Additionally, enhancing the explainability of AI-generated recommendations within FLoRA is essential for fostering transparency and trust among both learners and educators. Finally, dedicated training for teachers in the critical and effective use of GenAI tools will be crucial to maximize educational benefits while mitigating potential shortcomings.
In conclusion, by bridging the gap between automated AI-driven support and human agency, FLoRA contributes meaningfully to the development of HHAIRL research. Nevertheless, its effectiveness is contingent on thoughtful pedagogical integration, ongoing research, and proactive stakeholder engagement. Such a balanced approach will ensure that advances in AI not only empower learners to regulate their own learning but also enrich our collective understanding and practice of human-AI collaboration in education.
Declaration of Competing Interest
[…]
Appendix A Appendix - Other Features
@@ A.1 Other Tools
A.2 GenAI-based Scaffolding Triggering Logic
@@
This table describes the triggering logic of GenAI-based scaffolding tool.
Table 2: Learners’ learning conditions  (Li et al., 2025a)